OpenStudy (anonymous):

Show , lim x->0 ( sin 1/x) not equal to 0 use epsilon delta method

OpenStudy (anonymous):

it's lies in the interval -1,1 isn't ?

OpenStudy (anonymous):

nothing like that , prove by epsilon delta , thats all i know ! :p

OpenStudy (anonymous):

LOL!! Nice :))

OpenStudy (anonymous):

LOL ! :p

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=lim+x-%3E0+%28+sin+1%2Fx%29 seems like undefined

OpenStudy (anonymous):

is the ans there at this link ?

OpenStudy (experimentx):

it says that limit is undefined

OpenStudy (anonymous):

it's DNE!! :D

OpenStudy (anonymous):

but is it explained ?

OpenStudy (anonymous):

Nopes.

OpenStudy (anonymous):

lol ! I also know the ans , but the prob is explanation ! :p

OpenStudy (anonymous):

This has a somewhat rigorous proof

OpenStudy (anonymous):

This is a simple way to understand: http://mathforum.org/library/drmath/view/53369.html

OpenStudy (anonymous):

really ???

OpenStudy (anonymous):

Yes really!! :D

OpenStudy (anonymous):

thnksss !!!!! :->>

OpenStudy (anonymous):

Btw nice problem!

OpenStudy (anonymous):

I have many lyk these ! epsilon delta is really gettn on my nerves ! :p

OpenStudy (anonymous):

you are studying single variable or multivariate?

OpenStudy (anonymous):

both ! Actually in eco(h) , maths is subsi ! so its not core maths , but yea its maths , so bound to study ! :p

OpenStudy (anonymous):

lol, great :D :)) I wish I could help you more on epsilon delta proofs my current syllabus only requires to get the answer any how any ways ... so I tend to avoid rigorous stuffs.

OpenStudy (anonymous):

Lucky u ! :>

OpenStudy (anonymous):

lol, I have many many others things to study so not that :P

OpenStudy (anonymous):

Hint if you take the sequence \[ x_n = \frac{2}{(2n+1)\pi}\] The sequence \[ x_n\] converges to zero but \[ \sin\left ( \frac 1{x_n} \right)=\pm 1\]